# QCD Phase Diagram and the Finite Volume Fireball: A Model Study

### v1

@@ IV Number density and quark number susceptibility
The chiral condensate $\langle\bar{\psi}\psi\rangle$ and, equivalently, the constituent mass $M$ are good order parameters for tracking the chiral transition. They are, however, not directly connected to an experimental observable. The baryon number density ($n_{\scriptscriptstyle B}$) and susceptibilities of the baryon number operator are, on the other hand, related to experimental observables. In particular, they are of great interest in the large $\mu_{\scriptscriptstyle B}$ regime.
@@
$n_{q}=\frac{\partial P}{\partial\mu_{q}}=2N_{c}N_{f}\int\frac{d^{3}p}{(2\pi)^{3}}\left[n_{\scriptscriptstyle F}(E-\mu_{q};T)-n_{\scriptscriptstyle F}(E+\mu_{q};T)\right]\,,$ (8)
where $n_{\scriptscriptstyle F}$ is the Fermi distribution, $n_{\scriptscriptstyle F}(x;T)=\left(e^{\tfrac{x}{T}}+1\right)^{-1}$.
We start with an investigation of the effect of finite volume on $n_{q}$ in the crossover region. In the left panel of Fig. 6, we show the temperature dependence of $n_{q}$ at $\mu_{q}$=100 MeV. At small values of $\mu_{q}\lesssim 100$ MeV, the volume dependence of number density is small except in the crossover region (150 MeV $\leq\ T\ \leq$ 200 MeV). $n_{q}$ is small at small temperatures, and then rises and reaches its asymptotic behavior $\sim\mu_{q}T^{2}$ at high temperatures. In the crossover zone, we see some volume dependence: the rise of $n_{q}$ is milder at smaller volumes. This behavior gets further amplified as we go to higher values of $\mu_{q}$, where the transition becomes sharper. In the middle panel of Fig. 6 we show the results for $\mu_{q}$=250 MeV, still in the crossover region. Since the crossover is now sharper, $n_{q}$ increases much more sharply across the transition line at larger volume systems. Also above the transition the temperature is still not sufficiently high to show the asymptotic $\sim\mu T^{2}$ behavior. As we go to even higher values of $\mu_{q}$, we reach the first-order region, where $n_{q}$ shows a discontinuous jump across the transition. This is seen in the infinite volume case of the right panel of Fig. 6 ($\mu_{q}$=350 MeV). Also here the behavior of $n_{q}$ above the transition is quite different, as the $\mu_{q}^{4}$ term starts dominating.
@@
Figure 6: The quark number density ($n_{q}$) as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu_{q}$ of 100 MeV (left), 250 MeV (middle) and 350 MeV (right).
@@
The quark number density, and the nonlinear quark number susceptibilities which are higher-order derivatives of the free energy, are very interesting observables, especially for the critical point search. The nonlinear susceptibilities are defined as derivatives of the pressure ($P$) with respect to the quark chemical potential. Similar susceptibilities of other conserved quantities like charge or strangeness can also be constructed, but here we will concentrate on the quark number susceptibilities or equivalently, baryon number susceptibilities. The $n$th order susceptibility $\chi_{n}$ is defined as,
$\chi_{n}=\frac{\partial^{n}\,P}{\partial\,\mu_{q}^{n}}=\frac{\partial^{n-1}\,n_{q}}{\partial\,\mu_{q}^{n-1}}\,.$ (9)
The second-order susceptibility $\chi_{2}$ is a measure of the fluctuation of the baryon number. It shows a peak in the transition region. Near the critical region, the peak height becomes large, diverging at the critical point in the thermodynamic limit. In a finite volume system, the peak height scales with volume. In Fig. 7, we have shown $\chi_{2}$ at various values of $\mu_{q}$. Note the rapid increase in peak height and sharpened peak structure at the higher values of $\mu_{q}$.
Refer to caption Refer to caption Refer to caption
Figure 7: The quark number susceptibility $\chi_{2}$ as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu_{q}$ of 100 MeV (left), 250 MeV (middle) and 350 MeV (right).
@@
It has been suggested [27, 28] that the higher-order susceptibilities are a better probe for the critical region. In particular, $\chi_{3}$ [29] and $\chi_{4}$ [30, 31] show very interesting structures in the critical region, and these structures become smoother rapidly with decreasing volume. $\chi_{3}$ has a peak to the lower $\mu_{q}$ side of the transition line, then changes sign and shows a dip in the higher $\mu_{q}$ side. These peaks and dips become sharper as one approaches the critical point.
@@ V Summary of results and conclusion
The result of our investigation can be summarized by Fig. 5, which shows the shift of the transition line with the system volume. For the parameter sets used in our study, the NJL model has a first-order transition line at large $\mu_{q}$, ending in a critical point at $\left(T,\,\mu_{q}\right)\approx$ (60, 340) MeV. The actual location of the transition line as well as the endpoint is sensitive to the parameters chosen; the interest of this paper is on the sensitivity of the line on the system volume. Here the transition line has been tracked by the constituent quark mass $M$, or equivalently, $\langle\bar{\psi}\psi\rangle$. As the figure shows, in the crossover region, the shift of the transition line is mild: even for a system with $R$=5 fm, the shift is only a few MeV. The change of the number density with volume is shown in Fig. 6. As expected, the change of the number density across the crossover line becomes smoother as the volume is reduced.
The effect of the finite volume on the transition line is more interesting in the first-order region on the high $\mu_{q}$ side. The first-order region (indicated by the two-minima structure of the mean-field free energy; see Section 3) ends at a lower value of $T$ as the volume is reduced. The number density also shows very similar behavior. The volume dependence is particularly strong in the quark number susceptibilities of the baryon number in the critical region, as seen in Fig. 7. One would expect even larger effects for the higher-order susceptibilities. These results will be relevant for the experimental searches for the critical region, where the critical region is looked for in a small fireball.
@@ VI Acknowledgement
A.S. thank Prof. M. Chernodub for the helpful discussion. The computing resources of the Department of Theoretical Physics, TIFR, were used for this work. We would like to thank Ajay Salve and Kapil Ghadiali for their technical assistance. S.D. acknowledges the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002.
@@ Appendix A Calculational details
$\langle\bar{\psi}\psi\rangle_{\scriptscriptstyle\overline{MS}}=-\frac{N_{f}\,N_{c}\,M^{3}}{4\pi^{2}}\left(-1\,-\,\log\frac{\mu^{2}}{M^{2}}\right),$ (14)
where $\mu$ is the $\overline{MS}$ scale, and we have done the standard $\overline{MS}$ subtraction $\dfrac{1}{\epsilon}-\gamma+\log(4\pi)$ from the right hand side of Eq. (14). If we were to do the standard renormalization prescription of taking $\Lambda_{\rm\scriptscriptstyle UV}\to\infty$ in Eq. (13), then matching to Eq. (14) would require a subtraction of
$-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\Lambda_{\rm\scriptscriptstyle UV}^{2}\,+\,M^{2}\left(\log\frac{\mu^{2}}{\Lambda_{\rm\scriptscriptstyle UV}^{2}}\,+\gamma\right)\right]$
at this order. Of course, here $\Lambda_{\rm\scriptscriptstyle UV}$ is kept finite, and the result depends on the details of the regularization.
@@
$\langle\bar{\psi}\psi\rangle(T,\mu_{q})\,=\,-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\int_{\tau_{\rm\scriptscriptstyle UV}}^{\infty}\frac{d\tau}{\tau^{2}}\,e^{-M^{2}\tau}\;-\;\frac{1}{E_{p}}\left(n_{\scriptscriptstyle F}(E_{p}-\mu;T)\,+\,n_{\scriptscriptstyle F}(E_{p}+\mu;T)\right)\right],$ (16)
@@ Appendix B Comparison with other boundary conditions
The QCD phase diagram at finite volume has been studied in the literature, using the NJL model [33, 21] and other effective models like the PNJL [13, 14], the Polyakov quark-meson model [17, 34], etc., as well as using the Dyson-Schwinger equation [26] and the functional renormalization group [15]. As we have mentioned in Section II, most of the studies have used the periodic (PBC) or antiperiodic (APBC) boundary conditions. For a cubic box of volume $L^{3}$, this corresponds to momentum sum with spatial momenta $\dfrac{2n\pi}{L}$ and $\dfrac{(2n+1)\pi}{L}$, respectively. Also in some works [13, 14, 17], an infrared cutoff $\dfrac{\pi}{L}$ has been used in the momentum integral to mimic a system of dimensions $L$, but the geometry has not been specified. As we have mentioned in Section II, we do not think such boundary conditions mimic the finite size of the fireball. While a gross simplification, the MIT boundary condition used here captures the essential physics of the finite-size fireball.
@@
To see how much the results with the MIT boundary condition differ from the other boundary conditions, we have compared the size of the finite volume effects with the different boundary conditions, by looking at the constituent mass $M$. In Fig. 8, we have compared the results for the temperature dependence of $M$ for APBC and the IR cutoff for a box of dimension $L$=5 fm, with those of MIT boundary condition with $R$=5 fm.
Figure 8: Dressed quark mass ($M$) as a function of temperature ($T$) for different boundary conditions.
@@
As anticipated in Section II, the finite volume effect is significantly larger with the MIT boundary than the other boundary conditions. The APBC and the IR momentum cutoff show very little deviation from the infinite volume result already at $L$=5 fm. The PBC is expected to show even less finite volume effect than the APBC, as has been seen in lattice studies. In the chiral symmetry broken phase, the finite volume effect with the MIT boundary condition on a sphere is clearly more than that with APBC and IR cutoff, despite the volume of the sphere being about four times larger than those with a cubic geometry. In the high-temperature symmetry restored phase the boundary effect is substantially smaller for all the boundary conditions. Fig. 8 indicates, in our view, that the studies with the APBC and similar boundary conditions significantly underestimate the finite volume effect.
@@
While the APBC and PBC are not suitable for studying the finite volume fireball, they are excellent for numerical lattice studies, where one tries to get the infinite volume results from simulations on finite boxes. These boundary conditions are easy to implement on lattice, and show very little volume effect, making it easy to extrapolate to infinite volume from relatively small boxes. To check whether the mean-field NJL model qualitatively captures the volume dependence of the chiral condensate, we compared the results from lattice for $\langle\bar{\psi}\psi\rangle$ on small boxes with the NJL MF results. $\langle\bar{\psi}\psi\rangle$ is regularization dependent; in principle, one can convert them using results like those in Appendix A. Instead, we for simplicity compared the renormalization group invariant quantity:
@@
The comparison of such a study is shown in Fig. 9. This is a qualitative comparison where the lattice pion is somewhat heavier ($\approx$ 200 MeV). Also for the lattice, the $L\to\infty$ is replaced by the largest lattice we used. So a precise agreement is not expected. But as Fig. 9 shows, the deviation from the full theory at a very small volume is qualitatively captured in the NJL MF result. This gives us hope that the results we obtained for the MIT boundary condition, despite the many caveats, qualitatively capture the trend of the finite volume effect of a static fireball.

### v2

@@ IV Number density and quark number susceptibility
The chiral condensate $\langle\bar{\psi}\psi\rangle$ and, equivalently, the constituent mass $M$ are good order parameters for tracking the chiral transition. They are, however, not directly connected to an experimental observable. The baryon number density ($n_{\scriptscriptstyle B}$) and susceptibilities of the baryon number operator are, on the other hand, related to experimental observables. In particular, they are of great interest in the large $\mu_{\scriptscriptstyle B}$ regime, where one looks for the first order transition line and its endpoint.
@@
$n=\frac{\partial P}{\partial\mu}=2N_{c}N_{f}\int\frac{d^{3}p}{(2\pi)^{3}}\left[n_{\scriptscriptstyle F}(E-\mu;T)-n_{\scriptscriptstyle F}(E+\mu;T)\right]\,,$ (8)
where $n_{\scriptscriptstyle F}$ is the Fermi distribution, $n_{\scriptscriptstyle F}(x;T)=\left(e^{\tfrac{x}{T}}+1\right)^{-1}$. We start with an investigation of the effect of finite volume on $n$ in the crossover region. In the left panel of Fig. 6, we show the temperature dependence of $n$ at $\mu$=100 MeV. At small values of $\mu\lesssim 100$ MeV, the volume dependence of the number density is small except in the crossover region (150 MeV $\leq\ T\ \leq$ 200 MeV). $n$ is small at small temperatures, and then rises and reaches its asymptotic behavior $\sim\mu T^{2}$ at high temperatures. In the crossover zone, we see some volume dependence: the rise of $n$ is milder at smaller volumes. This behavior gets further amplified as we go to higher values of $\mu$, where the crossover becomes sharper. In the middle panel of Fig. 6 we show the results for $\mu$=250 MeV, still in the crossover region. Since the crossover is now sharper, $n$ increases much more sharply across the transition line at larger volume systems. As we enter the chirally symmetric phase, the temperature is still not sufficiently high to show the asymptotic $\sim\mu T^{2}$ behavior. As we go to even higher values of $\mu$, we reach the first-order region, where $n$ shows a discontinuous jump across the transition. This is seen in the infinite volume case of the right panel of Fig. 6 ($\mu$=350 MeV). The MIT sphere system has a smooth crossover at this value of $\mu$ for $R\leq$ 10 fm, while the 20 fm sphere is close to a transition, as is also seen in Fig. 5. The behavior of $n$ above the transition is quite different at such a large value of $\mu$, as the $\mu^{4}$ term starts dominating.
@@
Figure 6: The quark number density ($n$) as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu$ of 100 MeV (left), 250 MeV (middle) and 350 MeV (right).
@@
The quark number density, and the nonlinear quark number susceptibilities which are higher-order derivatives of the free energy, are very interesting observables, especially for the critical point search. The nonlinear susceptibilities are defined as derivatives of the pressure ($P$) with respect to the quark chemical potential. Similar susceptibilities of other conserved quantities like charge or strangeness can also be constructed, but here we will concentrate on the quark number susceptibilities or equivalently, baryon number susceptibilities. The $k$th order susceptibility $\chi_{k}$ is defined as,
$\chi_{k}=\frac{\partial^{k}\,P}{\partial\,\mu^{k}}=\frac{\partial^{k-1}\,n}{\partial\,\mu^{k-1}}\,.$ (9)
The behavior of $\chi_{k}$ can be inferred from the $\mu$ dependence of $n$, as sketched in the left panel of Fig. 7. At a fixed temperature, $n$ increases with $\mu$, with a rapid rise in the transition region. For a sufficiently high temperature $T>T_{c}$, the critical temperature, one has a crossover. $\chi_{2}$ therefore shows a moderate peak in this region. As the volume decreases, one expects a smoother crossover; therefore the peak height decreases. This behavior is seen in our calculated $\chi_{2}$ at different volumes; in the right panel of Fig. 7, we show the results at $\mu$ = 250 MeV, which is in the crossover region. Note also that the peak position has shifted slightly with volume. This is because of the shift of the crossover line, seen in Fig. 5.
Refer to caption           Refer to caption
Figure 7: (Left) Sketch of the $\mu$ dependence of the number density $n$ as well as the quark number susceptibilities. The dashed lines show the behavior in the crossover region, at a temperature $T>T_{c}$, the critical temperature. The solid lines correspond to a temperature $T<T_{c}$, where one has a first order transition. (Right) The quark number susceptibility $\chi_{2}$ as a function of $T$ on a system of different sizes (with MIT b.c.), at a quark chemical potential $\mu$=250 MeV, which is in the crossover region.
As one goes to lower temperatures, the crossover becomes steeper and eventually one approaches the critical point, where the $\chi_{2}$ peak is expected to diverge. In Fig. 8 we show $\chi_{2}$ in the low temperature region of the $\left(T,\,\mu\right)$ plane, for MIT spheres with $R$=10 and 5 fm. To obtain the curves, we calculated $n_{q}$ at a large number of closely spaced values of $\mu$, fitted a cubic spline, and took its derivative. We also checked that the values agreed with $\chi_{2}$ calculated directly at various points. As we see, the $\chi_{2}$ peak height increases as one approaches the critical point. As we discussed in Section III, the mean field effective potential shows a critical point also for the finite volume systems. This critical point was seen in Fig. 5 to move to lower temperatures with decreasing volume. The major volume effect seen in Fig. 8 is related to this shift in the critical point with volume. While the critical point in the infinite volume system appears at $T\approx$ 60 MeV, for the 10 fm sphere one only has a smooth crossover at this temperature. The critical point at this volume shows up at $T\approx$ 42 MeV, where the height of the $\chi_{2}$ peak is maximum. At lower temperatures, one has two separate phases at the transition: we calculate $\chi_{2}$ separately for the two phases, which show moderate values. The end point of the two-state transition happens at an even lower temperature for the 5 fm sphere, shifting the highest $\chi_{2}$ peak to $T\approx$ 20 MeV.
Figure 8: $\chi_{2}$ across the chiral transition line for $T\lesssim$ 60 MeV, which is the first order transition region for the infinite volume system. Shown are results for spheres of radii 5 and 10 fm.
@@
It has been suggested [33, 34] that the higher-order susceptibilities are a better probe for the critical region. In particular, $\chi_{3}$ [35] and $\chi_{4}$ [36, 37] show very interesting structures in the critical region, and these structures become smoother rapidly with decreasing volume. The behavior of $\chi_{3}$ across the crossover region has also been sketched in the left panel of Fig. 7. $\chi_{3}$ has a peak to the lower $\mu$ side of the transition line, then changes sign and shows a dip in the higher $\mu$ side. These peaks and dips become sharper as one approaches the critical point.
In the left panel of Fig. 9, we show the contour plot of $\chi_{3}$ calculated on a 10 fm sphere in the region $40\leq T\leq 60$ MeV, which is the region between the critical points on 10 fm and infinite volume systems. The qualitative behavior of $\chi_{3}$ seen in the figure is as sketched in Fig. 7. $\chi_{3}$ shows a peak as one approaches the transition region from the low $\mu$ side, then changes sign at the transition line and takes a large negative value just after the transition line, before becoming small again. The peak height and the valley depth increase as one approaches the critical region for this volume. Note that the structure for the 10 fm sphere is quite mild at $T$=60 MeV, which is close to the critical region for the infinite volume region. Similarly, for the 5 fm sphere, the peak and the valley are mild at $T$=40 MeV (compare the scales), but become steeper as one goes to lower temperatures, towards the endpoint of the discontinuity line for this volume.
Refer to caption           Refer to caption
Figure 9: Contour plots of $\chi_{3}/\mu$ near the endpoint of the first order line, for a sphere with a radius of 10 fm(left) and 5 fm (right) with MIT boundary condition. We have used Gaussian filtering to get rid of some spurious local structures which are numerical artifacts.
In the case of a conventional boundary condition like the antiperiodic one, the shift of the endpoint of first order line is insignificant for a box of size 5 fm or higher (see Appendix B), and the change in the magnitude of the $\chi_{2}$ peaks are due to the smoother nature of the transition at smaller volumes. As Fig. 8 and Fig. 9 show, here the major effect on the peak structure is due to the shift of the transition point. While our calculation is only at the mean field level, we believe this effect needs to be kept in mind when looking at the volume dependence of the baryon number cumulants (which are closely related to the susceptibilities).
@@ V Summary of results and conclusion
The result of our investigation of the phase diagram can be summarized by Fig. 5, which shows the shift of the transition line with the system volume. For the parameter sets used in our study, the NJL model has a first-order transition line at large $\mu$, ending in a critical point at $\left(T_{c},\mu_{c}\right)\approx$ (60, 340) MeV. The actual location of the transition line as well as the endpoint is sensitive to the parameters chosen; the interest of this paper is on the sensitivity of the line on the system volume. Here the transition line has been tracked by the constituent quark mass $M$, or equivalently, $\langle\bar{\psi}\psi\rangle$. As the figure shows, in the crossover region, the shift of the transition line is mild: even for a system with $R$=5 fm, the shift is only a few MeV. The volume dependence of the number density and its susceptibilities are discussed in Section IV. The volume dependence of the number density and its change across the transition line is shown in Fig. 6. As expected, the change of the number density across the crossover line becomes smoother as the volume is reduced.
The effect of the finite volume on the transition line is more interesting in the first-order region on the high $\mu$ side. The first-order region (indicated by the two-minima structure of the mean-field free energy; see Section 3) ends at a lower value of $T$ as the volume is reduced. The number density also shows very similar behavior. The volume dependence is particularly strong in the quark number susceptibilities of the baryon number in the critical region, as seen in Fig. 8. In particular, Fig. 8 shows that the major effect in the variation of the $\chi_{2}$ peak with volume is due to the shift of the endpoint of the first order line. One expects a sharp peak at the transition region near this point. As the location of the endpoint shifts with volume, the region of large $\chi_{2}$ peaks also shifts. This effect seems to be more significant than the expected sharpening of the peak heights with volume. Similarly, $\chi_{3}$ shows a very interesting behavior across the transition region, with sharp peaks and valleys around the transition line. These structures are seen in Fig. 9 to show an even stronger shift with volume, due to the shift in the location of the critical point. One would expect even larger effects for the higher-order susceptibilities. These results will be relevant for the experimental searches for the critical region, where the critical region is looked for in a small fireball.
@@ VI Acknowledgement
A.S. thank Prof. M. Chernodub for a helpful discussion. The computing resources of the Department of Theoretical Physics, TIFR, were used for this work. We would like to thank Ajay Salve and Kapil Ghadiali for their technical assistance. S.D. acknowledges the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002.
@@ Appendix A Calculational details
$\langle\bar{\psi}\psi\rangle_{\scriptscriptstyle\overline{MS}}=-\frac{N_{f}\,N_{c}\,M^{3}}{4\pi^{2}}\left(-1\,-\,\log\frac{\overline{\mu}^{2}}{M^{2}}\right),$ (14)
where $\overline{\mu}$ is the $\overline{MS}$ scale, and we have done the standard $\overline{MS}$ subtraction $\dfrac{1}{\epsilon}-\gamma+\log(4\pi)$ from the right hand side of Eq. (14). If we were to do the standard renormalization prescription of taking $\Lambda_{\rm\scriptscriptstyle UV}\to\infty$ in Eq. (13), then matching to Eq. (14) would require a subtraction of
$-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\Lambda_{\rm\scriptscriptstyle UV}^{2}\,+\,M^{2}\left(\log\frac{\overline{\mu}^{2}}{\Lambda_{\rm\scriptscriptstyle UV}^{2}}\,+\gamma\right)\right]$
at this order. Of course, we keep $\Lambda_{\rm\scriptscriptstyle UV}$ finite, and the results depend on the details of the regularization.
@@
$\langle\bar{\psi}\psi\rangle(T,\mu)\,=\,-\frac{N_{f}\,N_{c}\,M}{4\pi^{2}}\left[\int_{\tau_{\rm\scriptscriptstyle UV}}^{\infty}\frac{d\tau}{\tau^{2}}\,e^{-M^{2}\tau}\;-\;\frac{1}{E_{p}}\left(n_{\scriptscriptstyle F}(E_{p}-\mu;T)\,+\,n_{\scriptscriptstyle F}(E_{p}+\mu;T)\right)\right],$ (16)
@@ Appendix B Comparison with other boundary conditions
The QCD phase diagram at finite volume has been studied in the literature, using the NJL model [39, 24] and other effective models like the PNJL [16, 17], the linear sigma and the Polyakov quark-meson model [15, 20, 40], etc., as well as using the Dyson-Schwinger equation [31] and the functional renormalization group [18]. As we have mentioned in Section II, most of the studies have used the periodic (PBC) or the antiperiodic (APBC) boundary condition. For a cubic box of volume $L^{3}$, they correspond to momentum sum with spatial momenta $\dfrac{2n\pi}{L}$ and $\dfrac{(2n+1)\pi}{L}$, respectively. Also in some works [16, 17, 20], an infrared cutoff $\dfrac{\pi}{L}$ has been used in the momentum integral to mimic a system of dimensions $L$, but the geometry has not been specified. As we have mentioned in Section II, we do not think such boundary conditions mimic the finite size of the fireball. The MIT boundary condition used here captures the essential physics, though in a very simplified form, of the finite-size fireball.
@@
To see how much the results with the MIT boundary condition differ from the other boundary conditions, we have compared the size of the finite volume effects with the different boundary conditions, by looking at the constituent mass $M$. In the left panel of Fig. 10, we have compared the results for the temperature dependence of $M$ at $\mu=0$ for APBC and the IR cutoff for a box of dimension $L$=5 fm, with those obtained for a sphere with $R$=5 fm with MIT boundary condition.
@@
As anticipated in Section II, the finite volume effect is significantly larger with the MIT boundary than what is seen with the other boundary conditions. The APBC and the IR momentum cutoff show very little deviation from the infinite volume result even at $L$=5 fm. The PBC is expected to show even less finite volume effect than the APBC, as has been seen in lattice studies. In the chiral symmetry broken phase, the finite volume effect with the MIT boundary condition on a sphere is clearly more than that with APBC and IR cutoff, despite the volume of the sphere being about four times larger than those with a cubic geometry. In the high-temperature symmetry restored phase the boundary effect is substantially smaller for all the boundary conditions.
Refer to caption       Refer to caption
Figure 10: (Left) Constituent quark mass ($M$) as a function of temperature ($T$) at $\mu=0$ for different boundary conditions. (Right) The transition/crossover line in the $\left(T,\,\mu\right)$ plane for a small system with MIT and antiperiodic boundary conditions, compared with the infinite volume line. In both these figures the MIT boundary condition is imposed on a sphere with radius $R$ =5 fm, while the other boundary conditions use a cubic box with side $L$= 5 fm.
In the right panel of the figure, we have compared the phase transition line in the $\left(T,\,\mu\right)$ plane for an infinite volume system, along with those obtained for a finite volume box of dimension 5 fm with APBC and a sphere of radius 5 fm with MIT boundary condition. The interpretation of the line is similar to that used in Section III and in Fig. 5: it is the line obtained from the $M$ dependence of the mean field effective potential. The antiperiodic box shows very little volume dependence even at 5 fm: the crossover line is very close to the infinite volume one, and even the coexistence region in $M$ starts to appear at a very similar values of $\left(T,\,\mu\right)$  [^2] For the linear sigma model, a significant shift of the transition line, and shift of the endpoint of the first order line towards lower temperatures, have been reported for a 5 fm box with APBC and PBC [15].. At lower temperatures, the “first order line” (in the mean field effective potential sense) shows slight deviation from the infinite volume case, but still within the coexistence region. On the other hand, as discussed in Section III, the MIT boundary condition leads to a substantial change: the crossover line is pushed to higher values of $\mu$ at all temperatures. Also the coexistence region is shrunk significantly at this volume. Fig. 10 supports our intuition that the studies with the APBC and similar boundary conditions significantly underestimate the finite volume effect.
@@
While the APBC and the PBC are not suitable for studying the finite volume fireball, they are excellent for numerical lattice studies, where one tries to get the infinite volume results from simulations on finite boxes. These boundary conditions are easy to implement on lattice, and show very little volume effect, making it easy to extrapolate to infinite volume from relatively small boxes. To check whether the mean-field NJL model qualitatively captures the volume dependence of the chiral condensate, we compared the results from lattice for $\langle\bar{\psi}\psi\rangle$ on small boxes with the NJL MF results. $\langle\bar{\psi}\psi\rangle$ is regularization dependent; in principle, one can use results like those in Appendix A to compare the results from different regularizations. Instead, we for simplicity compared the renormalization group invariant quantity:
@@
The result of such a comparison is shown in Fig. 11. This is a qualitative comparison where the lattice pion is somewhat heavier ($\approx$ 200 MeV). Also for the lattice, the $L\to\infty$ is replaced by the largest lattice we used. So a precise agreement is not expected. But as Fig. 11 shows, the deviation from the full theory at a very small volume is qualitatively captured in the NJL MF result. This gives us hope that the results we obtained for the MIT boundary condition, despite the many caveats, qualitatively capture the trend of the finite volume effect of a static fireball.